OpenStudy (anonymous):

If Bob gave 2/5 of his pennies to Beverly and Beverly gave 2/5 of the pennies she received to Susan then Susan gave 2/5 of the pennies she received to Donald. (Donald received 64 pennies) How many pennies did Bob initially have?

OpenStudy (pfenn1):

Let x = pennies Bob initially had. How many pennies did Bob give to Beverly?

OpenStudy (pfenn1):

*Let x = number of pennies Bob initially had

OpenStudy (anonymous):

what would be the equation I would use ?

OpenStudy (pfenn1):

Bob gave 2/5 of his pennies to Beverly therefore Beverly got (2/5)x pennies where x = number of pennies Bob initially had. Now Beverly gave 2/5 of the number of pennies she had to Susan therefor Susan got (2/5)(2/5)x pennies. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and donald recieved 64 pennies which was 2/5^3 of the original pennies

OpenStudy (pfenn1):

Right!

OpenStudy (anonymous):

I know this but how wold I come to the result? What equation would I use?

OpenStudy (pfenn1):

The number of pennies Donald received is (2/5)^3 of the original pennies Bob started with. Can you write an equation that expresses that idea?

OpenStudy (anonymous):

tried but no success which is why I posted here to find an equation that would work

OpenStudy (anonymous):

so can you give me an equation?

OpenStudy (pfenn1):

number of pennies Donald received = 64 x= number of pennies Bob started with so the equation is\[64=\left( \frac 25 \right)^3x\]

OpenStudy (pfenn1):

Rearrange and solve for x.

OpenStudy (lgbasallote):

wouldnt it be neater to simplify it already?\[\large 64 = \frac{8x}{125}\]

OpenStudy (anonymous):

@crzymom2u are you following?

OpenStudy (anonymous):

got the answer is 1000

OpenStudy (pfenn1):

Yes!